# Broken Expectations: The Effects of Modelling Assumptions on the Inferred Dark Matter Distribution in the Milky Way’s Satellites

### v1

@@ Abstract
The spherical Jeans equation is commonly used to infer dark matter distributions in dwarf spheroidal satellites of the Milky Way to constrain the nature of dark matter. One of its assumptions is that of dynamical equilibrium while the dwarfs are under the influence of Galactic tides. We carry out tailored simulations of Carina, Draco, Fornax, Sculptor and Ursa Minor and analyse the accuracy of dark matter density profiles and annihilation rates (J-factors) recovered with the Jeans equation, using pyGravSphere. We find that tides do not significantly affect the quality of density profile inference when using a bound sample of stars in the 5 simulated dwarfs; however, pyGravSphere tends to underestimate the inner densities of dwarf galaxies, which, together with tidal mass loss, leads to an inference of flatter density slopes, although all of our dwarfs have cuspy Navarro-Frenk-White haloes. The recovered J-factors are generally underestimated. While the difference with the true J-factor is small, the error bars are often underestimated and must be at least of order 0.1 dex to encompass the true J-factor. We also test the accuracy of the Wolf et al. 2010 mass estimator and find that it can be sensitive to orbital stage and eccentricity. Still, for our sample of dwarf galaxies, the estimates agree with the truth within 10 per cent. Consistency of our simulated dwarfs with the mass-concentration relation in CDM requires a light Milky Way, or limited action of tides, which may be in tension with a “tidal stirring” origin of dwarf spheroidals.
@@
Within the standard -Cold Dark Matter (CDM) model of cosmology (Davis et al., 1985), dark matter makes up per cent of the matter density in our Universe, yet its nature remains unknown (Planck Collaboration et al., 2020). In the last two decades, dwarf galaxies have become of great interest in dark matter searches (Strigari, 2018). Their high dynamical mass-to-light ratios (McConnachie, 2012) imply that the internal kinematics of these objects are strongly dark matter dominated, making it possible to study the nature of dark matter phenomenologically through its dynamical effect on the baryonic matter in dwarf galaxies (Łokas, 2002). Moreover, since the Milky Way is currently known to host 66 satellite dwarf galaxies and more potential candidates (DES Collaboration, 2015; Drlica-Wagner et al., 2020; Homma et al., 2024; Delve Collaboration, 2025), their relative proximity implies high fluxes of the by-products of dark matter annihilation or decay, making it possible to constrain the particle physics nature of dark matter with gamma and X-ray instruments such as Fermi (Fermi-LAT Collaboration, 2015; McDaniel et al., 2024), the Cherenkov Telescope Array (Lefranc et al., 2016; Saturni et al., 2023), XMM-Newton and eRosita (Saeedi & Sasaki, 2022; Nucita et al., 2013; Ando et al., 2021).
@@
A key property relevant for dark matter studies is its density distribution near the centres of galaxies. In particular, some -body and galaxy formation simulations in CDM predict that the dark matter density should be increasing towards the halo centres (Navarro et al., 1996b; Springel et al., 2008a; Fattahi et al., 2016a; Bose et al., 2019), forming a “cusp” (). Other simulations (Tollet et al., 2016; Fitts et al., 2017; Orkney et al., 2021; Jackson et al., 2024), as well as some measurements of rotation curves in nearby dwarf galaxies, seem to suggest that the dark matter density is constant at galaxy centres (Flores & Primack, 1994; Moore, 1994; Oh et al., 2015; Adams et al., 2014), forming a “core” (). This has become known as the “core-cusp problem” (de Blok, 2010), which remains a challenge to the standard model of cosmology. Two main solutions have been proposed to this problem: one is that bursty supernovae feedback in shallow dwarf galaxy potentials is responsible for lowering the dark matter density over time (Navarro et al., 1996a; Read & Gilmore, 2005; Mashchenko et al., 2008; Pontzen & Governato, 2012; Brooks & Zolotov, 2014; Read et al., 2016), the other is that dark matter is not cold, or can interact non-gravitationally, producing a core at galaxy centres (Spergel & Steinhardt, 2000). The latter explanation has motivated many studies in alternative dark matter models (Macciò et al., 2012; Elbert et al., 2015; Schive et al., 2014; Chan et al., 2022). The shortcoming of some of these models, however, is that they tend to produce dark matter cores along the full range of dark matter halo masses and in all halos, whereas latest rotation curve measurements imply a variety of dark matter distribution – cusped, cored and something in-between (de Naray et al., 2010; Santos-Santos et al., 2020). Some models are additionally inconsistent with sizes and density distributions inferred for nearby dwarfs (Correa, 2021; Dalal & Kravtsov, 2022; Zimmermann et al., 2025). As such, the “core-cusp problem” in recent years has transformed into a “rotation curve diversity problem” (Oman et al., 2015; Relatores et al., 2019), requiring a solution that can produce a variety of dark matter density distributions, with self-interacting dark matter, featuring core formation and gravothermal core collapse that can be accelerated by the presence of tidal forces (Kahlhoefer et al., 2019; Nishikawa et al., 2020), becoming a popular candidate (Kamada et al., 2017; Tulin & Yu, 2018; Roberts et al., 2024).
@@
The difficulty in obtaining a convincing detection of a dark matter core through the modelling of stellar kinematics is three-fold. Firstly, the kinematic and photometric samples of stars in the Milky Way dwarf spheroidals are typically small, owing partially to the already small stellar masses of these galaxies (typically < 10). Chang & Necib (2021) and Guerra et al. (2023) have shown that samples of stars are required for a convincing detection of a dark matter core. Secondly, the kinematic data is typically limited to the line-of-sight velocities, leading to a so-called mass-anisotropy degeneracy in the popular Jeans dynamical models, meaning that the uncertainty on the transverse motions of the stars translates to the uncertainty on the enclosed mass profile. Even with the proper motion data becoming available with Gaia (Gaia Collaboration et al., 2021; Li et al., 2021), most dwarfs of interest are too far away to allow for sufficiently small proper motion errors (Strigari et al., 2018); however, some works have combined the Gaia proper motions with Hubble to reduce these errors (Massari et al., 2020; Vitral et al., 2024). Finally, dynamical models (particularly those based on the Jeans equation) typically make a number of assumptions which are known to be broken in Milky Way’s dwarf satellites: spherical symmetry (McConnachie, 2012), dispersion-dominated internal motions (Martínez-García et al., 2021) and dynamical equilibrium (Newberg et al., 2003; del Pino et al., 2015). This introduces a degree of scepticism in the interpretation of the results of dynamical models which do not account for these effects and has led several works to evaluate how cored or cuspy dark matter haloes are based on their densities below the half-light radius compared to CDM predictions – a quantity that can be evaluated more reliably than the innermost log-slope of the dark matter density distribution (Read et al., 2019; Gregory et al., 2019; Collins et al., 2021; Charles et al., 2023; De Leo et al., 2024).
@@
A largely unexplored source of uncertainty in the literature is the effect of Milky Way’s tidal forces on the inferred dark matter distribution in the inner regions of dwarf galaxies, particularly for several satellites in a consistent Milky Way and cosmological framework; however, for individual satellites, several works have found limited impact by tides and accurate mass profile recovery given effective removal of unbound stars (Klimentowski et al., 2007, 2009; Serra et al., 2010; Ural et al., 2015; Battaglia et al., 2015; Read et al., 2018; Iorio et al., 2019; De Leo et al., 2024). Cosmological galaxy formation simulations in CDM have made it possible to test the accuracy of modern dynamical models on simulated dwarfs. Genina et al. (2020) have tested the pyGravSphere method on Fornax-like dwarfs in the APOSTLE (Sawala et al., 2016; Fattahi et al., 2016b) simulations (A Project for Simulating the Local Environment) and found no significant effect on the inner profiles and a significant effect on the outermost profiles, primarily caused by the limited priors to model the steepening of dark matter slope in the outer halo caused by tides (Peñarrubia et al., 2009). Wang et al. (2022) have tested Jeans anisotropic modelling (Cappellari et al., 2013) on Sagittarius-like dwarfs with the AURIGA simulations (Grand et al., 2017) and found that their enclosed masses and inner density profiles were typically underestimated due to contraction motion caused by tidal effects. Nguyen et al. (2025) tested a Graph Neural Posterior Estimator on a sample of dwarfs from the FIRE-2 Latte simulations (Wetzel et al., 2016) and have identified trends of the method’s effectiveness with the orbital pericentres, how many pericentres have been completed and the current position with respect to the host galaxy, with stronger tidal effects leading to an overestimation of the dwarf galaxies’ masses and maximum circular velocities. Nevertheless, it is unclear whether these results are applicable to Milky Way’s dwarfs, as it is unlikely that any given simulation would reproduce both the observed properties of the dwarf galaxies and their present-day orbits, thus making them unconvincing analogues of Milky Way’s satellites.
@@
In this work, we circumvent these issues by producing a suite of tailored simulations of Carina, Draco, Fornax, Sculptor and Ursa Minor dwarf spheroidals in two disky Milky Way potentials from Bovy (2015). Our simulated dwarf galaxies follow orbits in these potentials as derived from Gaia EDR3 proper motions (Battaglia et al., 2022) and, at , have stellar masses, half-light radii and line-of-sight velocity dispersions consistent with current observations. Our dwarfs are all set up to have spherical Navarro-Frenk-White (NFW) dark matter density profiles initially (Navarro et al., 1996b) and stellar components following a Plummer distribution (Plummer, 1911). In future work, we will aim to replicate this experiment with cored dark matter density profiles. For now, we aim to establish whether current observations of Milky Way’s classical dwarfs can be replicated in a CDM cosmology where cores do not form. We then focus exclusively on the effect of Milky Way’s tidal forces on the ability of the spherical Jeans models to recover underlying dark matter density distributions of dwarf satellites and the accuracy of their recovered J-factors.
@@ 2.1.1 Milky Way Potential
For this work, we use and modify the standard Milky Way potential contained within the galpy Python library (Bovy, 2015), dubbed MWPotential2014. The models used in this potential were fitted with various data sources, the details of which are further elaborated in Bovy & Rix (2013) and Bovy (2015). The potential consists of an NFW dark matter halo, a power-law bulge potential and a Miyamoto-Nagai disk potential (Miyamoto & Nagai, 1975). For the purposes of this work, we modify MWPotential2014 as done in Battaglia et al. (2022) by increasing the virial mass, , of the dark matter halo to 1.6, while also using the original virial mass of 0.8 so that we have a ‘heavy’ and ‘light’ Milky Way potential, respectively, which are in the upper and lower range of virial masses estimated for the Milky Way in the literature (Wang et al., 2020). We note, however, that this is not a fully realistic choice of the Milky Way potential for several reasons: it does not take the time evolution of the gravitational potential due to the Milky Way’s assembly (Correa et al., 2015; Evans et al., 2020) into account, nor the adiabatic contraction of the potential due to the accumulation of gas in the halo centre (Cautun et al., 2020), or asymmetries due to the rotating Galactic bar (Binney & Tremaine, 2008) or massive satellites (Garavito-Camargo et al., 2019), although it is now well established that the Large Magellanic Cloud is likely to have strongly influenced the orbits of Milky Way’s satellites (Patel et al., 2020). It should also be noted that the Bovy (2015) potential has a very low dark matter fraction in the centre, which is incompatible with cosmological simulations in CDM and may make it difficult to sustain a stable disk (Lovell et al., 2018). We also do not take dynamical friction into account in this work, as we found that it only results in minor deviations of the dwarf spheroidal orbits (the greatest effect is found in Fornax, where the orbit is delayed by 0.2 Gyr with the last apocentre offset by no more than 10 kpc, compared to the prediction of the Chandrasekhar approximation (Chandrasekhar, 1943) as implemented in galpy111We note that there is dynamical self-friction present in our simulations, which also partially causes a decrease in dSph orbital apocentres.). Although it is likely that our simulated satellite orbits do not follow the true ones, there is equally much uncertainty in the Milky Way mass (Wang et al., 2020), the mass of the LMC (Erkal et al., 2019; Watkins et al., 2024), whether the satellites had a past association with LMC (Santos-Santos et al., 2021) and how strongly the satellites are affected by dynamical friction due to their mass and the mass distribution in the Milky Way (Santistevan et al., 2023). These uncertainties would introduce an enormous set of parameters to be considered, and we thus focus on satellite orbits in a static Milky Way potential without including these additional unknowns. Taking the results of Patel et al. (2020) and Battaglia et al. (2022) at face value, this most likely means that our simulated satellites are affected by tides more strongly for ‘heavy’ and ‘light’ Milky Way potentials, as interaction with LMC favours a late infall for most of our satellites, although the cumulative effect of tidal forces to the present day may be somewhere between the ‘light’ and ‘heavy’ cases, particularly if the dwarfs had undergone tidal pre-processing as part of group infall (Genina et al., 2022; Jahn et al., 2022; Callingham & Helmi, 2025).

### v2

@@ Abstract
The spherical Jeans equation is commonly used to infer dark matter distributions in dwarf spheroidal satellites of the Milky Way to constrain the nature of dark matter. One of its assumptions is that of dynamical equilibrium while the dwarfs are under the influence of Galactic tides. We carry out tailored simulations of Carina, Draco, Fornax, Sculptor and Ursa Minor and test the accuracy of dark matter density profiles and annihilation rates (J-factors) recovered with the Jeans analysis code pyGravSphere. We find that tides do not significantly affect the quality of density profile inference; however, pyGravSphere tends to underestimate the inner densities of dwarf galaxies, which, together with tidal mass loss, leads to an inference of flatter density slopes, although all of our dwarfs have cuspy Navarro-Frenk-White haloes. This is because the default broken power-law model is unable to describe the outer halo density profile. The recovered J-factors are generally underestimated. While the difference with the true J-factor is small, the error bars are also often underestimated. We also test the accuracy of the Wolf et al. 2010 mass estimator and find that it can be sensitive to orbital stage and eccentricity. Still, for our sample of dwarf galaxies, the estimates agree with the truth within 10 per cent. Consistency of our simulated dwarfs with the mass-concentration relation in CDM requires a light Milky Way, or limited action of tides, which may be in tension with a “tidal stirring” origin of dwarf spheroidals.
@@
Within the standard -Cold Dark Matter (CDM) model of cosmology (Davis et al., 1985), dark matter makes up per cent of the matter density in our Universe, yet its nature remains unknown (Planck Collaboration et al., 2020). In the last two decades, dwarf galaxies have become of great interest in dark matter searches (Baltz et al. 1999; Evans et al. 2004; Bergström & Hooper 2006; Sánchez-Conde et al. 2007; Strigari et al. 2008; see Strigari 2018 for a review). Their high dynamical mass-to-light ratios (Mateo, 1998) imply that the internal kinematics of these objects are strongly dark matter dominated (Battaglia et al., 2013; Walker, 2013; Battaglia & Nipoti, 2022), making it possible to study the nature of dark matter phenomenologically through its dynamical effect on the baryonic matter in dwarf galaxies (Wilkinson et al., 2002; Łokas, 2002). Moreover, since the Milky Way is currently known to host 60 confirmed and candidate satellite dwarf galaxies (DES Collaboration, 2015; Drlica-Wagner et al., 2020; Homma et al., 2024; Delve Collaboration, 2025; Pace, 2025), their relative proximity implies high fluxes of the by-products of dark matter annihilation or decay, making it possible to constrain the particle physics nature of dark matter with gamma and X-ray instruments such as Fermi (Fermi-LAT Collaboration, 2015; McDaniel et al., 2024), the Cherenkov Telescope Array (Lefranc et al., 2016; Saturni et al., 2024), XMM-Newton and eRosita (Saeedi & Sasaki, 2022; Nucita et al., 2013; Ando et al., 2021).
@@
A key property relevant for dark matter studies is its density distribution near the centres of galaxies. In particular, some -body dark-matter-only (Navarro et al., 1996b; Springel et al., 2008a) and galaxy formation simulations in CDM (Fattahi et al., 2016a; Bose et al., 2019) predict that the dark matter density should be increasing towards the halo centres, forming a “cusp” (). Other simulations111See Benítez-Llambay et al. (2019) on the likely source of difference between simulation predictions. (Tollet et al., 2016; Fitts et al., 2017; Orkney et al., 2021; Jackson et al., 2024), as well as some measurements of rotation curves in nearby dwarf galaxies, seem to suggest that the dark matter density is constant at galaxy centres (Flores & Primack, 1994; Moore, 1994; Oh et al., 2015; Adams et al., 2014), forming a “core” (). This has become known as the “core-cusp problem” (de Blok, 2010), which remains a challenge to the standard model of cosmology (see Sales et al. (2022) for a review). Two main solutions have been proposed to this problem: one is that bursty supernovae feedback in shallow dwarf galaxy potentials is responsible for lowering the dark matter density over time (Navarro et al., 1996a; Read & Gilmore, 2005; Mashchenko et al., 2008; Pontzen & Governato, 2012; Brooks & Zolotov, 2014; Read et al., 2016), the other is that dark matter is not cold, or can interact non-gravitationally, producing a core at galaxy centres (Spergel & Steinhardt, 2000). The latter explanation has motivated many studies in alternative dark matter models (Macciò et al., 2012; Elbert et al., 2015; Schive et al., 2014; Chan et al., 2022). The shortcoming of some of these models, however, is that they tend to produce dark matter cores along the full range of dark matter halo masses and in all halos, whereas latest rotation curve measurements imply a variety of dark matter distribution – cusped, cored and something in-between (de Naray et al., 2010; Santos-Santos et al., 2020). Some models are additionally inconsistent with sizes and density distributions inferred for nearby dwarfs (Correa, 2021; Dalal & Kravtsov, 2022; Zimmermann et al., 2025). As such, the “core-cusp problem” in recent years has transformed into a “rotation curve diversity problem” (Oman et al., 2015; Relatores et al., 2019), requiring a solution that can produce a variety of dark matter density distributions, with self-interacting dark matter, featuring core formation and gravothermal core collapse that can be accelerated by the presence of tidal forces (Kahlhoefer et al., 2019; Nishikawa et al., 2020), becoming a popular candidate (Kamada et al., 2017; Tulin & Yu, 2018; Roberts et al., 2025).
@@
The difficulty in obtaining a convincing detection of a dark matter core through the modelling of stellar kinematics is three-fold. Firstly, the kinematic data are typically limited to the line-of-sight velocities, leading to a so-called mass-anisotropy degeneracy in the popular Jeans dynamical models, meaning that the uncertainty on the transverse motions of the stars translates to the uncertainty on the enclosed mass profile. Even with the proper motion data becoming available with Gaia (Gaia Collaboration et al., 2021; Li et al., 2021), most dwarfs of interest are too far away to allow for sufficiently small proper motion errors (Strigari et al., 2018); however, some works have combined the Gaia proper motions with the Hubble Space Telescope data to reduce these errors (Massari et al., 2020), while some utilised the longer-baseline HST data, reaching deeper magnitude limits than Gaia (Vitral et al., 2024a, b). Secondly, the kinematic and photometric samples of stars in the Milky Way dwarf spheroidals are typically small, owing partially to the already small stellar masses of these galaxies (typically < 10). Chang & Necib (2021) and Guerra et al. (2023) have shown that samples of kinematic tracers are required for a convincing detection of a dark matter core222Although we note that these studies consider only a single available tracer population and only use the constraints of the second moment of the velocity distribution., although this may depend on core size. Finally, dynamical models (particularly those based on the Jeans equation) typically make a number of assumptions which are known to, or might to some extent, be broken in Milky Way’s dwarf satellites, such as spherical symmetry (McConnachie, 2012) and dynamical equilibrium (Newberg et al., 2003; del Pino et al., 2015). This introduces a degree of scepticism in the interpretation of the results of dynamical models which do not account for these effects and has led several works to evaluate how cored or cuspy dark matter haloes are based on their densities below the half-light radius compared to CDM predictions – a quantity that can be evaluated more reliably than the innermost log-slope of the dark matter density distribution (Read et al., 2019; Gregory et al., 2019; Collins et al., 2021; Charles et al., 2023; De Leo et al., 2024). Although rotation can be accounted for in Jeans and other types of models (e.g. Zhu et al. 2016), its effects when modelling Milky Way’s dwarf spheroidals are commonly assumed to be negligible, with observational studies confirming that rotation is typically insignificant compared to the velocity dispersion (Wheeler et al., 2017; Martínez-García et al., 2021). The parameterization of the dynamical model and modelling technique can also lead to significant differences in the inferred dark matter distribution (Read et al., 2021).
@@
A largely unexplored source of uncertainty in the literature is the effect of Milky Way’s tidal forces on the inferred dark matter distribution in the inner regions of dwarf galaxies, particularly for several satellites in a consistent Milky Way and cosmological framework; however, for individual satellites, several works have found limited impact by tides and accurate mass profile recovery given effective removal of unbound stars (Klimentowski et al., 2007, 2009; Serra et al., 2010; Ural et al., 2015; Battaglia et al., 2015; Iorio et al., 2019; De Leo et al., 2024). Cosmological galaxy formation simulations in CDM have made it possible to test the accuracy of modern dynamical models on simulated dwarfs. Genina et al. (2020) have tested the pyGravSphere method on Fornax-like dwarfs in the APOSTLE (Sawala et al., 2016; Fattahi et al., 2016b) simulations (A Project for Simulating the Local Environment) and found no significant effect on the inner profiles and a significant effect on the outermost profiles, primarily caused by the limited priors to model the steepening of dark matter slope in the outer halo caused by tides (Peñarrubia et al., 2009). Wang et al. (2022) have tested Jeans anisotropic modelling (Cappellari et al., 2013) on Sagittarius-like dwarfs with the AURIGA simulations (Grand et al., 2017) and found that their enclosed masses and inner density profiles were typically underestimated due to contraction motion caused by tidal effects. Nguyen et al. (2025) tested a Graph Neural Posterior Estimator on a sample of dwarfs from the FIRE-2 Latte simulations (Wetzel et al., 2016) and have identified trends of the method’s effectiveness with the orbital pericentres, how many pericentres have been completed and the current position with respect to the host galaxy, with stronger tidal effects leading to an overestimation of the dwarf galaxies’ masses and maximum circular velocities. Nevertheless, it is unclear whether these results are applicable to Milky Way’s dwarfs, as it is unlikely that any given simulation would reproduce both the observed properties of the dwarf galaxies and their present-day orbits, thus making them unconvincing analogues of Milky Way’s satellites.
@@
As in the past works by Battaglia et al. (2015) and Iorio et al. (2019), who studied tidal effects on Fornax and Sculptor dSphs, we circumvent these issues by producing a suite of tailored simulations of Carina, Draco, Fornax, Sculptor and Ursa Minor in two disky Milky Way potentials from Bovy (2015). Our simulated dwarf galaxies follow orbits in these potentials as derived from Gaia EDR3 proper motions (Battaglia et al., 2022) and, at , have stellar masses, half-light radii and line-of-sight velocity dispersions consistent with current observations. Our dwarfs are all set up to have spherical Navarro-Frenk-White (NFW) dark matter density profiles initially (Navarro et al., 1996b) and stellar components following a Plummer distribution (Plummer, 1911). The dark matter content of the dwarfs is resolved with particles to avoid the effects of artificial disruption (van den Bosch & Ogiya, 2018). In future work, we will aim to replicate this experiment with cored dark matter density profiles. For now, we aim to establish whether current observations of Milky Way’s classical dwarfs can be replicated in a CDM cosmology where cores do not form. We then focus exclusively on the effect of Milky Way’s tidal forces on the ability of spherical Jeans models to recover underlying dark matter density distributions of dwarf satellites and the accuracy of their recovered J-factors. We do not consider the effects of unbound stars in contaminating the kinematic sample and focus instead on the deformation of the dwarf spheroidal system due to tides and the tidal heating of particle orbits that need not necessarily leave the stars unbound. We make our sample of simulated dwarf spheroidals publicly available for tests of alternative dynamical models with varying kinematic sample selections.
@@ 2.1.1 Milky Way Potential
For this work, we use and modify the standard Milky Way potential contained within the galpy Python library (Bovy, 2015), dubbed MWPotential2014. The models used in this potential were fitted with various data sources, the details of which are further elaborated in Bovy & Rix (2013) and Bovy (2015). The potential consists of an NFW dark matter halo, a power-law bulge potential and a Miyamoto-Nagai disk potential (Miyamoto & Nagai, 1975). For the purposes of this work, we modify MWPotential2014 as done in Battaglia et al. (2022) by increasing the virial mass, , of the dark matter halo to 1.6, while also using the original virial mass of 0.8 so that we have a ‘heavy’ and ‘light’ Milky Way potential, respectively, which are in the upper and lower range of virial masses estimated for the Milky Way in the literature (Wang et al., 2020a). We note, however, that this is not a fully realistic choice of the Milky Way potential for several reasons: it does not take the time evolution of the gravitational potential due to the Milky Way’s assembly (Correa et al., 2015; Evans et al., 2020) into account, nor the adiabatic contraction of the potential due to the accumulation of gas in the halo centre (Cautun et al., 2020), or asymmetries due to the rotating Galactic bar (Binney et al., 1997) or massive satellites (Garavito-Camargo et al., 2019), although it is now well established that the Large Magellanic Cloud is likely to have strongly influenced the orbits of Milky Way’s satellites (Patel et al., 2020; Battaglia et al., 2022; Correa Magnus & Vasiliev, 2022). It should also be noted that the Bovy (2015) potential has a very low dark matter fraction in the centre, which is incompatible with cosmological simulations in CDM and may make it difficult to sustain a stable disk (Lovell et al., 2018). We also do not take dynamical friction into account in this work, as we found that it only results in minor deviations of the dwarf spheroidal orbits (the greatest effect is found in Fornax, where the orbit is delayed by 0.2 Gyr with the last apocentre offset by no more than 10 kpc, compared to the prediction of the Chandrasekhar approximation (Chandrasekhar, 1943) as implemented in galpy333We note that there is dynamical self-friction present in our simulations, which also partially causes a decrease in dSph orbital apocentres.). Although it is likely that our simulated satellite orbits do not follow the true ones, there is equally much uncertainty in the Milky Way mass (Wang et al., 2020a), the mass of the LMC (Erkal et al., 2019; Watkins et al., 2024), whether the satellites had a past association with LMC (Santos-Santos et al., 2021) and how strongly the satellites are affected by dynamical friction due to their mass and the mass distribution in the Milky Way (Santistevan et al., 2023). These uncertainties would introduce an enormous set of parameters to be considered, and we thus focus on satellite orbits in a static Milky Way potential without including these additional unknowns. Taking the results of Patel et al. (2020) and Battaglia et al. (2022) at face value, this most likely means that our simulated satellites are affected by tides more strongly for ‘heavy’ and ‘light’ Milky Way potentials, as interaction with LMC favours a late infall for most of our satellites, although the cumulative effect of tidal forces to the present day may be somewhere between the ‘light’ and ‘heavy’ cases, particularly if the dwarfs had undergone tidal pre-processing as part of group infall (Genina et al., 2022; Jahn et al., 2022; Callingham & Helmi, 2025).